PLEASE HELP! One pipe can fill a swimming pool in 5 hours less than another. Together they fill the swimming pool in 5 hours. How long would it take each pipe to fill the tank alone?

Answer by AnlytcPhil(1806)

PLEASE HELP! One pipe can fill a swimming pool in 5 hours less than another.
Together they fill the swimming pool in 5 hours. How long would it take each
pipe to fill the tank alone?

Make this chart:

             D = deeds done =           R =              T =   
           No. of pools filled    Rate in pools/hr   Time in hours
Fast pipe                                                      
Slow pipe              
Both together                                             

Let the time for the slow pipe = x
The time for fast pipe is x-5

Fill in the times.

             D = deeds done =           R =              T =   
           No. of pools filled    Rate in pools/hr   Time in hours
Fast pipe                                               x-5    
Slow pipe                                                x    
Both together                                            5

Now use the fact that in all three cases the number of pools filled is 1.

             D = deeds done =           R =              T =   
           No. of pools filled    Rate in pools/hr   Time in hours
Fast pipe           1                                   x-5    
Slow pipe           1                                    x    
Both together       1                                    5

Now use R = D/T to fill in the rates

             D = deeds done =           R =              T =   
           No. of pools filled    Rate in pools/hr   Time in hours
Fast pipe           1                 1/(x-5)           x-5    
Slow pipe           1                   1/x              x    
Both together       1                   1/5              5

Now we form our equation from

Rate of fast pipe + Rate of slow pipe = Their combined rate

(That reasoning is similar to the fact that when a boat is traveling
downstream we add the rate of the boat to the rate of the river to
find their combined rate.)

1/(x-5) + 1/x = 1/5

Can you solve that?  It requires the quadratic formula
                      _
Answers:  x = (15 ± 5Ö5)/2 approximately =  13.09 hrs and 1.91 hrs. for the
time of the slow pipe.

We discard the 1.91 hours because that would give a negative number for the
hours it takes the fast pipe to fill the pool.

So the answers are, approximately: It takes the slow pipe 13.09 hours and 
the fast pipe 5 hours less, or 8.09 hours. 

Edwin
AnlytcPhil@aol.com